OpenStudy (anonymous):

The function s(V) = \[\sqrt[3]{x}\] describes the side length, in units, of a cube with a volume of V cubic units. Jason wants to build a cube with a minimum of 64 cubic centimeters. What is a reasonable range for s, the side length, in centimeters, of Jason’s cube?

OpenStudy (anonymous):

\[s \ge 0\]\[s \ge 4\]\[s \ge 8\]\[s \ge 16\]

OpenStudy (dayakar):

\[\sqrt[3]{x}=64\] since volume of square equal to 64

OpenStudy (dayakar):

can u find x value

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

no, I don't think I ever learned how to do that

OpenStudy (dayakar):

\[\sqrt[3]{x }=x ^{\frac{ 1 }{ 3 }}\] x^1/3 = 4^3 { 64 = 4*4*4 = 4^3]

OpenStudy (dayakar):

i will explain it again \[s(v) =\sqrt[3]{x}\] \[s(64) = \sqrt[3]{64}\]

OpenStudy (anonymous):

so B, \[s \ge 4\]

OpenStudy (dayakar):

exactly